# MirrorGuard: Adaptive Defense Against Jailbreaks via Entropy-Guided Mirror Crafting

### v1

@@ Abstract
Defending large language models (LLMs) against jailbreak attacks is crucial for ensuring their safe deployment. Existing defense strategies generally rely on predefined static criteria to differentiate between harmful and benign prompts. However, such rigid rules are incapable of accommodating the inherent complexity and dynamic nature of real jailbreak attacks. In this paper, we propose a novel concept of “mirror” to enable dynamic and adaptive defense. A mirror refers to a dynamically generated prompt that mirrors the syntactic structure of the input while ensuring semantic safety. The personalized discrepancies between the input prompts and their corresponding mirrors serve as the guiding principles for defense. A new defense paradigm, MirrorGuard, is further proposed to detect and calibrate risky inputs based on such mirrors. An entropy-based detection metric, Relative Input Uncertainty (RIU), is integrated into MirrorGuard to quantify the discrepancies between input prompts and mirrors. MirrorGuard is evaluated on several popular datasets, demonstrating state-of-the-art defense performance while maintaining general effectiveness.
MirrorGuard: Adaptive Defense Against Jailbreaks via Entropy-Guided Mirror Crafting
@@ 1 Introduction
Securing LLMs has emerged as a critical challenge due to their widespread deployment in various essential domains (OpenAI, 2023; Zhu et al., 2024). Operating within open scenarios, LLMs are inherently susceptible to jailbreak attacks, which exploit vulnerabilities to circumvent LLM safety protocols, thereby compelling models to generate harmful outputs (Liu et al., 2024; Chen et al., 2024).
Defense strategies against jailbreak attacks can be broadly categorized into prefill-level (Xie et al., 2023; Phute et al., 2023) and response-level approaches (Phute et al., 2023). Prefill-level defenses enhance the robustness of initial prompts or input structures, using methods like protective prefixes and perplexity-based classification to filter malicious inputs (Zheng et al., 2024; Alon and Kamfonas, 2023). In contrast, response-level strategies focus on evaluating model-generated outputs, applying safety filters to mitigate potential harm (Phute et al., 2023; Xu et al., 2024).
@@
Figure 1: The differences between static discrimination-based defense and our proposed method.
The nucleus in defending LLMs against jailbreak attacks lies in accurately distinguishing harmful (“bad”) input prompts from harmless (“good”) ones, with the precise definition of “good” and “bad” serving as the cornerstone. Existing approaches primarily rely on heuristically defined rules to classify input prompts as harmful or benign (Yi et al., 2024). Prefill-level defenses generally employ pre-defined rules (e.g., keyword lists  (Robey et al., 2023)) or fixed input features (e.g., perplexity thresholds (Alon and Kamfonas, 2023)) to identify malicious prompts. Similarly, response-level defenses leverage post-generation filters, such as harmful content detectors, or pre-established evaluation metrics, including LLM-based evaluations, to flag risky outputs (Wang et al., 2024). However, these static, pre-defined criteria are inherently limited in addressing the nuanced and sophisticated nature of adversarial inputs (Dong et al., 2024). As depicted in Figure 1(a), jailbreak attack prompts manifest in a variety of forms and styles, thereby rendering pre-defined static criteria overly simplistic and rigid. Frameworks that rely on static discrimination between harmful and benign prompts cannot accommodate the inherent complexity and variability of jailbreak attacks, resulting in suboptimal performance in sophisticated real-world scenarios (Huang et al., 2024).
Different from previous approaches relying on static discrimination, we aim to learn dynamic distinction rules that adaptively evaluate the risk of input prompts. Inspired by the concept of Reflective Equilibrium (Rawls, 1971), which achieves consistency through iterative adjustments between judgments and principles, we propose first identifying the input prompts that deviate from the “mirror” and then iteratively adjusting these deviations. The mirror is defined as a dynamically generated prompt that mirrors the syntactic structure of the input while ensuring semantic safety. As illustrated in Figure 1(b), each input prompt is paired with a mirror, and LLMs generate responses for both. If the response to the input significantly deviates from that of its mirror, it is likely to be risky. By incorporating the discrepancy between the responses of the input and its mirror, the discrimination standard becomes dynamically adaptive to each input, as the mirror is generated in a personalized manner. Furthermore, the mirror also serves as a guide for iteratively adjusting the input by minimizing the deviations between the responses of the vanilla input and its mirror.
While mirrors provide a dynamic reference for distinguishing harmful from harmless prompts, their generation and utilization pose significant challenges. First, an effective method is needed to dynamically and accurately generate a suitable mirror that represents the harmless counterpart of the input prompt. Second, a robust metric is essential to quantify the discrepancies between the input and its mirror. Third, an iterative strategy is required to reduce discrepancies between the LLM’s response and the mirror, ultimately ensuring the generation of a safe output.
@@
In this paper, we propose a novel defense method, MirrorGuard, to tackle the aforementioned challenges. Our motivation lies in generating mirrors to serve as dynamic benchmarks for detecting jailbreak attack prompts and guiding the target LLM to produce safe outputs. MirrorGuard consists of three primary modules: the mirror maker, the mirror selector, and the entropy defender. The mirror maker generates desirable mirrors based on the input prompt, implemented as an instruction fine-tuned model. Once the mirrors are generated, the mirror selector identifies the optimal mirror by comparing its syntactic similarity and semantic harmlessness to the input prompt. Finally, the entropy defender iteratively refines the input prompt via mirror comparison, aligning its Relative Input Uncertainty (RIU) value with that of the mirror. Experimental results on popular datasets demonstrate the superiority of our proposal. Our contributions are summarized as follows:
@@
  • To the best of our knowledge, we are the first to investigate the novel problem of dynamic and adaptive defense against jailbreak attacks, based on the innovative concept of “mirror”.
@@
  • We propose MirrorGuard, a novel defense model that dynamically generates mirrors to adaptively detect attacks. An iterative alignment strategy is proposed to minimize deviations between inputs and their mirrors.
  • We conduct extensive experiments on popular benchmarks and various jailbreak attack methods, demonstrating that our approach outperforms baselines in reducing attack success rates while maintaining overall performance.
@@
Figure 2: The overview of the proposed MirrorGuard model, including the mirror maker, the mirror selector, and the entropy defender via mirror comparison.
@@ 2 Preliminary
Entropy is a foundational concept in information theory (Pimentel et al., 2021). As a measure of uncertainty in a LLM’s output, entropy reflects the model’s confidence of its information content (Araabi et al., 2024). Given a random variable $X$ with the set $[x_{0},\dots,x_{n}]$ of possible outcomes, the Shannon entropy (Shannon, 1948) is defined as follows:
@@ 3 Methodology
Figure 2 (a) illustrates the joint framework of the MirrorGuard paradigm. Given a input jailbreak attack prompt $p$, MirrorGuard firstly generates a set of candidate mirrors {$l_{1}$, $l_{2}$,…, $l_{n}$} via the Mirror Maker. Then, Mirror Selector selects the most satisfied mirror $l_{i}^{\prime}$ which has the same linguistic characters with the input $p$. $l_{i}^{\prime}$ is expected to compare the behavioral inconsistencies in the sentence understanding of the LLM. Finally, Entropy Defender via Mirror Comparison calculates RIU between current input and mirrors to identify the jailbreak attack prompt, and guides the LLM toward generating more confident and reliable responses through the multiple-query guidance process.
3.1 Mirror Maker
The objective of mirror maker is to produce mirrors that share the similar linguistic features with the input jailbreak attack prompt. An effective mirror is expected to satisfy three constraints. (1) Length constraint: the mirror is expected to have a token count similar to that of the input prompt, allowing for a fair comparison between prompt pairs (Steindl et al., 2024). (2) Syntax Constraint: the mirror needs to follow syntactic rules consistent with the input prompt to control variables in the LLM’s comprehension of the input (Zhu et al., 2024). (3) Sentiment constraint: the mirror is also intended to have a non-negative sensitive polarity, serving as a safe and standard reference (Huang et al., 2024).
Constrained Instruction Tuning. In light of the stochastic nature of LLM’s outputs and their reliance on general-purpose training, it is intractable for existing models to directly generate mirrors due to the lack of explicit control mechanisms. Drawing inspiration from prior work (Zhou et al., 2023), we employ fine-tuning guided by natural language instructions and demonstrations to construct an instruction-tuned model.
@@ 3.1 Mirror Maker
After the process of data synthesis, the model is fine-tuned by using a combination of instruction tuning (Wei et al., 2022) and meta-in-context learning (Min et al., 2022). For each training example, the natural language instructions for the constraints are prepended with 5 demonstrations, which consist of constraint–output pairs. These demonstrations are either of the same constraint type or a composition of multiple constraints. The model is then fine-tuned by using maximum likelihood estimation and teacher forcing technology to ensure training stability and constraint compliance (Zhou et al., 2023).
@@ 3.2 Mirror Selector
To evaluate and filter the candidate mirrors, an LLM-based classifier (e.g., GPT-4o) is employed. This classifier assesses each candidate mirror against the three constraints and assigns judgment tags, such as “true” or “false” to indicate whether the criteria are satisfied. Only the mirrors that simultaneously satisfy all three criteria are considered suitable. Among these suitable mirrors, the first five are selected as the final choices to ensure robustness and effectiveness. The details of mirror selector process are provided in Appendix B.
@@ 3.3 Entropy Defender via Mirror Comparison
Following the mirror maker and mirror selector, a set of desirable mirrors is generated. However, the integration of these mirrors for effective defense remains obscure. We further propose the entropy defender to leverage the discrepancies between input prompts and their mirrors to dynamically assess and mitigate the risks.
@@ 3.3.2 RIU-based Defense
From Equation (6), when the input prompt is benign, the numerator and denominator tend to be similar (i.e., $RIU_{attention}$ close to 1.0), as both quantify the deviations between two safe yet syntactically similar sentences. If $RIU_{attention}$ exceeds a predefined threshold $\sigma$ (close to 1.0), then the LLM is classified as harmless, and the model proceeds with response generation. However, if $RIU_{attention}$ falls below the threshold, the input is flagged as potentially harmful, triggering a multiple-query guidance process. This process iteratively refines the input prompt through strategies such as simplification (Steindl et al., 2024)(e.g., “Please simplify the current sentence”) and multiple queries (Hu et al., 2024) until $RIU_{attention}$ exceeds the threshold. The algorithm for MirrorGuard is detailed in Appendix C.
@@ 4.1 Experimental Settings
Target LLMs. The defense method is evaluated on two widely used open-source LLMs: Llama2-7b-chat (Touvron et al., 2023) and Vicuna-7b (Chiang et al., 2023). Following previous work (Zhou et al., 2023), T0-11b is utilized as the base model of instruction-tuned model because it has previously been instruction-tuned on many NLP tasks.
Attacks. To address the effectiveness of MirrorGuard, six representative attacks are selected to be compared: GCG (Zou et al., 2023), AutoDAN (Liu et al., 2024), PAIR (Chao et al., 2023), TAP (Mehrotra et al., 2023), DrAttack (Li et al., 2024) and BaitAttack (Pu et al., 2024).
Baselines. Following the previous work (Xu et al., 2024), five SOTA defense mechanisms are considered as baselines, including detection-based (Perplexity Filter (Alon and Kamfonas, 2023)) and mitigation-based methods (ICD (Wei et al., 2023), Paraphrase (Jain et al., 2023), Self-Reminder (Xie et al., 2023) and SafeDecoding (Xu et al., 2024)).
@@
AdvBench HEx-PHI GCG AutoDAN PAIR TAP DrAttack BaitAttack
Llama2 No Defense 0.000 0.002 0.382 0.575 0.174 0.187 0.511 0.654 0.414
Perplexity Filter 0.000 0.002 0.000 0.575 0.174 0.187 0.511 0.654 0.350
ICD 0.000 0.000 0.000 0.000 0.147 0.156 0.231 0.255 0.132
Paraphrase 0.002 0.003 0.004 0.000 0.124 0.151 0.275 0.298 0.142
Self-Reminder 0.000 0.000 0.000 0.000 0.139 0.147 0.301 0.342 0.155
SafeDecoding 0.000 0.001 0.004 0.000 0.004 0.005 0.009 0.011 0.006
MirrorGuard(Ours) 0.000 0.001 0.000 0.000 0.002 0.002 0.006 0.009 0.003
Vicuna No defense 0.008 0.172 0.910 0.847 0.815 0.823 0.854 0.979 0.871
Perplexity Filter 0.008 0.154 0.000 0.847 0.815 0.823 0.854 0.979 0.720
ICD 0.000 0.006 0.711 0.805 0.537 0.544 0.832 0.891 0.720
Paraphrase 0.143 0.237 0.225 0.714 0.679 0.680 0.811 0.887 0.666
Self-Reminder 0.000 0.008 0.428 0.702 0.485 0.486 0.825 0.854 0.630
SafeDecoding 0.000 0.001 0.000 0.000 0.008 0.009 0.013 0.025 0.009
MirrorGuard (Ours) 0.000 0.001 0.000 0.000 0.002 0.007 0.010 0.012 0.005
@@
MirrorGuard(Ours) 1.043 $\times$ 1.075 $\times$
Table 2: This table summarizes ATGR of MirrorGuard and the baseline defense approaches.
@@
MirrorGuard(Ours) 68.6% 0.443 91.2% 0.535
@@
Metrics. Following previous work Xu et al. (2024), the attack success rate (ASR) is used to measure the effectiveness of the defense methods, which is defined as: $ASR=\frac{\#\text{successful prompts}}{\#\text{total prompts}}$. The success of a jailbreak attack is evaluated by GPT-4o (Qi et al., 2023). The Average Token Generation Time Ratio (ATGR) is used to assess the time cost of all defense methods (Xu et al., 2024). It is defined as $ATGR=\frac{T_{\text{defense}}}{T_{\text{non-defense}}}$, where $T_{\text{defense}}$ indicates the average token generation time when defenses are implemented, and $T_{\text{non-defense}}$ indicates the average token generation time when defenses are not implemented. Moreover, the WinRate and Rouge-L scores are used to evaluate the general performance of LLMs in dealing with harmless tasks (Jiang et al., 2024).
Implementation Details. The details of implementation settings are given in Appendix B.

### v2

# Rui Pu1, Chaozhuo Li1, Rui Ha1, Litian Zhang2, Lirong Qiu1, Xi Zhang1

@@ Abstract
Defending large language models (LLMs) against jailbreak attacks is crucial for ensuring their safe deployment. Existing defense strategies typically rely on predefined static criteria to differentiate between harmful and benign prompts. However, such rigid rules fail to accommodate the inherent complexity and dynamic nature of real-world jailbreak attacks. In this paper, we focus on the novel challenge of universal defense against diverse jailbreaks. We propose a new concept “mirror”, which is a dynamically generated prompt that reflects the syntactic structure of the input while ensuring semantic safety. The discrepancies between input prompts and their corresponding mirrors serve as guiding principles for defense. A novel defense model, MirrorShield, is further proposed to detect and calibrate risky inputs based on the crafted mirrors. Evaluated on multiple benchmark datasets and compared against ten state-of-the-art attack methods, MirrorShield demonstrates superior defense performance and promising generalization capabilities.
[Uncaptioned image] MirrorShield: Towards Universal Defense Against Jailbreaks via Entropy-Guided Mirror Crafting
@@ 1 Introduction
Securing LLMs has emerged as a critical challenge due to their widespread deployment in essential domains (OpenAI, 2023; Zhu et al., 2024). Operating in open scenarios, LLMs are inherently susceptible to jailbreak attacks, which exploit vulnerabilities to circumvent safety protocols and induce harmful outputs (Liu et al., 2024b; Chen et al., 2024). Current defense strategies can be broadly categorized into prefill-level methods that block harmful prompts during input (Zheng et al., 2024; Alon and Kamfonas, 2023; Xie et al., 2023), and response-level approaches which focus on post-generation filtering or output evaluation (Phute et al., 2023; Xu et al., 2024; Wang et al., 2025).
@@
Figure 1: The differences between static discrimination-based defense and our proposed dynamic method.
Despite their progress, existing defense approaches often target only specific types of attack methods, making them prone to failure in practice as jailbreak attacks grow increasingly sophisticated and diverse (Dong et al., 2024a). As depicted in Figure 1(a), jailbreak attacks manifest in a variety of forms and styles, in which obfuscated jailbreak attacks using indirect or multilingual phrasing have been proven to easily circumvent keyword-based defense models (Pingua et al., 2024).
The root cause of this dilemma stems from the static and fixed definition of what constitutes harmful prompts (Dong et al., 2024b). Existing defense models predominantly rely on heuristically defined criteria to locate harmful prompts, including predefined keyword lists (Robey et al., 2023), perplexity-based thresholds (Alon and Kamfonas, 2023), and post-generation filters (Wang et al., 2024). However, such static, pre-defined criteria are overly simplistic and rigid in adapting to the evolving and sophisticated nature of adversarial inputs (Dong et al., 2024a). For example, keyword-based filters inherently rely on superficial lexical patterns, reflecting a static view of harmfulness that fails to capture the dynamic attack patterns. As a result, defense models that depend on static discrimination criteria achieve suboptimal performance in sophisticated real-world scenarios (Huang et al., 2024).
In this paper, we aim to enhance the universality of jailbreak defenses by shifting from fixed static discrimination to dynamic relative judgment. Instead of relying on absolute markers of harmful prompts, we propose a novel paradigm of contrastive comparison based on the theory of relativity (Baratz, 1978). Our insight lies in the idea that judgments of good or bad lack standalone meaning and derive significance only through comparison with alternative viewpoints. Since directly evaluating a single prompt’s harmfulness is challenging, we propose comparing it to its “mirror” reflection. The mirror is defined as a dynamically generated prompt that mirrors the syntactic structure of the input while ensuring semantic safety. As shown in Figure 1(b), each input prompt is paired with its mirrored counterpart, and the LLM generates responses for both. Prior research reveals that LLMs differentiate between harmful and harmless prompts through internal representations (e.g., attention patterns or hidden states) (Zhou et al., 2024b), making the activation discrepancy triggered by inputs and their mirrors a valuable risk indicator. The mirror is dynamically generated to align with the input’s specific characteristics, ensuring versatile applicability against diverse attack formats. Furthermore, the mirror acts as a security reference to detect subtle semantic distinctions within input-mirror pairs, thereby effectively countering sophisticated and varied jailbreak prompts.
While mirrors serve as dynamic references for identifying harmful prompts, their generation and utilization present three key challenges. First, an effective method is required to dynamically generate accurate mirrors that represent harmless counterparts of input prompts. Second, robust metrics are essential for quantifying discrepancies between inputs and their mirrors. Third, although the discrepancy within input-mirror pairs indicates risks, relying solely on this metric may lead to false positives by falsely rejecting benign inputs. Thus, an ideal defense model should smoothly generate safe responses rather than rigidly rejecting risky inputs.
@@
In this paper, we propose a novel defense method, MirrorShield, to tackle the aforementioned challenges. Our motivation lies in creating mirrors to serve as dynamic benchmarks for detecting jailbreak attack prompts and guiding the target LLM to produce safe outputs, analogous to the Mirror Shield in the game Zelda. MirrorShield consists of three primary modules: the mirror generator, the mirror selector, and the entropy defender. The mirror generator utilizes the constrained instruction tuning to ensure the mirrors are harmless counterparts of inputs. Once the mirrors are generated, the mirror selector identifies the optimal mirror by comparing its syntactic similarity and semantic harmlessness to the input prompt. Finally, the entropy defender iteratively refines the input through comparison with its mirror, guiding the model to generate a safe and appropriate response by gradually narrowing the gap between their discrepancy. Experimental results on popular datasets demonstrate the superiority of our proposal. Our contributions are summarized as follows:
@@
  • We investigate the novel problem of universal defense against diverse emerging jailbreak attack methods through the dynamic and adaptive innovative concept of “mirror”.
@@
  • We propose MirrorShield, a novel defense model that dynamically generates mirrors to adaptively detect attacks. An iterative alignment strategy is introduced to steer the model toward generating safer and more appropriate responses with the guidance of mirror.
  • We conduct extensive experiments on popular benchmarks and various jailbreak attacks, demonstrating the effectiveness and generality of our proposal.
@@
Figure 2: The overview of the proposed MirrorShield model, including the mirror generator, the mirror selector, and the entropy defender via mirror comparison.
@@ 2 Preliminary
As a measure of uncertainty in a LLM’s output, entropy reflects the model’s confidence of its information content (Pimentel et al., 2021; Araabi et al., 2024). Given a random variable $X$ with the set $[x_{0},\dots,x_{n}]$ of possible outcomes, the Shannon entropy (Shannon, 1948) is defined as:
@@ 3 Methodology
Figure 2 (a) illustrates the joint framework of the MirrorShield paradigm. Given a input jailbreak attack prompt $p$, MirrorShield firstly generates a set of candidate mirrors {$l_{1}$, $l_{2}$,…, $l_{n}$} via the Mirror Generator. Then, Mirror Selector selects the most satisfied mirror $l_{i}^{\prime}$ which has the same linguistic characters with the input $p$. $l_{i}^{\prime}$ is expected to compare the behavioral inconsistencies in the sentence understanding of the LLM. Finally, Entropy Defender via Mirror Comparison calculates RIU between current input and mirrors to identify the jailbreak attack prompt, guiding the LLM toward generating more confident and reliable responses through the multiple-query guidance process.
3.1 Mirror Generator
The objective of mirror generator is to produce mirrors that share the similar linguistic features with the input jailbreak attack prompt. An effective mirror is expected to satisfy three constraints. (1) Length constraint: the mirror is expected to have a token count similar to that of the input prompt, allowing for a fair comparison between prompt pairs (Steindl et al., 2024). (2) Syntax Constraint: the mirror needs to follow syntactic rules consistent with the input prompt to control variables in the LLM’s comprehension of the input (Zhu et al., 2024). (3) Sentiment constraint: the mirror is also intended to have a non-negative sensitive polarity, serving as a safe and standard reference (Huang et al., 2024). Although the mirror follows the syntactic structure of the original prompt, it does not preserve the original intent, as it is generated by a controllable text model rather than a conversational language model. The generation process is guided only by structural patterns, while the content is regenerated under sentiment control. As a result, maintaining non-negative sentiment serves as an effective way to ensure semantic safety.
Constrained Instruction Tuning. In light of the stochastic nature of LLM’s outputs and their reliance on general-purpose training, it is intractable for existing models to directly generate mirrors due to the lack of explicit control mechanisms. Inspired by prior work (Zhou et al., 2023), we employ fine-tuning guided by natural language instructions to construct an instruction-tuned model.
@@ 3.1 Mirror Generator
After the process of data synthesis, the model is fine-tuned by using a combination of instruction tuning (Wei et al., 2022) and meta-in-context learning (Min et al., 2022). For each training example, the instructions for the constraints are prepended with 5 demonstrations, which consist of constraint–output pairs. These demonstrations are either of the same constraint type or a composition of multiple constraints. The model is then fine-tuned by using maximum likelihood estimation and teacher forcing technology to ensure constraint compliance (Zhou et al., 2023).
@@ 3.2 Mirror Selector
To evaluate and filter the candidate mirrors, an LLM-based classifier (e.g., GPT-4o) is employed. This classifier assesses each candidate mirror against the three constraints and assigns judgment tags, such as “true” or “false” to indicate whether the criteria are satisfied. Only the mirrors that simultaneously satisfy all three criteria are considered suitable. Among these suitable mirrors, the first five are selected as the final choices to ensure robustness and effectiveness. The details of mirror selector process are provided in Appendix C.
@@ 3.3 Entropy Defender via Mirror Comparison
Following the mirror generator and mirror selector, a set of desirable mirrors is generated. However, the integration of these mirrors for effective defense remains obscure. We further propose the entropy defender to leverage the discrepancies between input prompts and their mirrors to dynamically assess and mitigate the risks.
@@ 3.3.2 RIU-based Defense
From Equation (6), when the input prompt is benign, the numerator and denominator tend to be similar (i.e., $RIU_{attention}$ close to 1.0), as both quantify the deviations between two safe yet syntactically similar sentences. If $RIU_{attention}$ exceeds a predefined threshold $\sigma$ (close to 1.0), then the LLM is classified as harmless, and the model proceeds with response generation. However, if $RIU_{attention}$ falls below the threshold, the input is flagged as potentially harmful, triggering a multiple-query guidance process. This process iteratively refines the input prompt through strategies such as simplification (Steindl et al., 2024)(e.g., “Please simplify the current sentence”) and multiple queries (Hu et al., 2024) until $RIU_{attention}$ exceeds the threshold. The algorithm for MirrorShield is detailed in Appendix D.
@@ 4.1 Experimental Settings
Target LLMs. The defense method is evaluated on three widely used open-source LLMs: Llama2-7b-chat (Touvron et al., 2023), Vicuna-13b-v1.5 (Chiang et al., 2023) and Mistral-7b. Following previous work (Zhou et al., 2023), T0-11b is utilized as the base model of instruction-tuned model because it has been instruction-tuned on many NLP tasks.
Attacks. To address the effectiveness of MirrorShield, ten representative attacks are selected to be compared: GCG (Zou et al., 2023), AutoDAN (Liu et al., 2024b), SAP30 (Deng et al., 2023), DeepInception (Li et al., 2023), GPTFuzzer-Template (Yu et al., 2023), PAIR (Chao et al., 2023), TAP (Mehrotra et al., 2023), DrAttack (Li et al., 2024), BaitAttack (Pu et al., 2024) and DRA (Liu et al., 2024a).
Baselines. Following the previous work (Xu et al., 2024), five SOTA defense mechanisms are considered as baselines, including detection-based (Perplexity Filter (Alon and Kamfonas, 2023)) and mitigation-based methods (ICD (Wei et al., 2023), Paraphrase (Jain et al., 2023), Self-Reminder (Xie et al., 2023), SafeDecoding (Xu et al., 2024) and Layer-AdvPatcher (Ouyang et al., 2025)).
@@
AdvBench HEx-PHI GCG AutoDAN SAP30 Deep Template PAIR TAP DrAttack BaitAttack DRA
Llama2 No Defense 0.000 0.002 0.382 0.275 0.000 0.181 0.215 0.174 0.187 0.511 0.654 0.568 0.315
Perplexity Filter 0.000 0.002 0.000 0.275 0.000 0.181 0.215 0.174 0.187 0.511 0.654 0.568 0.277
ICD 0.000 0.000 0.000 0.000 0.000 0.000 0.154 0.147 0.156 0.231 0.255 0.248 0.119
Paraphrase 0.002 0.003 0.004 0.000 0.000 0.082 0.049 0.124 0.151 0.275 0.298 0.210 0.119
Self-Reminder 0.000 0.000 0.000 0.000 0.000 0.038 0.197 0.139 0.147 0.201 0.242 0.288 0.125
SafeDecoding 0.000 0.001 0.004 0.000 0.000 0.016 0.108 0.004 0.005 0.009 0.011 0.037 0.019
Layer-AdvPatcher 0.000 0.003 0.142 0.208 0.000 0.219 0.177 0.165 0.170 0.419 0.502 0.535 0.254
MirrorShield(Ours) 0.000 0.001 0.000 0.000 0.000 0.013 0.015 0.002 0.002 0.006 0.009 0.015 0.006
Vicuna No Defense 0.080 0.172 0.895 0.817 0.825 0.458 0.498 0.609 0.637 0.810 0.974 0.819 0.734
Perplexity Filter 0.080 0.154 0.000 0.817 0.825 0.458 0.498 0.609 0.637 0.810 0.974 0.819 0.645
ICD 0.000 0.006 0.684 0.805 0.456 0.421 0.461 0.410 0.428 0.535 0.562 0.519 0.528
Paraphrase 0.143 0.237 0.245 0.714 0.549 0.397 0.359 0.331 0.344 0.632 0.656 0.598 0.483
Self-Reminder 0.000 0.008 0.403 0.702 0.428 0.452 0.455 0.429 0.435 0.413 0.439 0.453 0.461
SafeDecoding 0.000 0.018 0.071 0.064 0.143 0.118 0.075 0.080 0.090 0.033 0.052 0.102 0.083
Layer-AdvPatcher 0.106 0.237 0.846 0.759 0.598 0.435 0.475 0.517 0.533 0.789 0.921 0.765 0.664
MirrorShield(Ours) 0.000 0.011 0.000 0.000 0.060 0.048 0.063 0.020 0.050 0.010 0.028 0.076 0.036
Mistral No Defense 0.061 0.154 0.826 0.768 0.789 0.416 0.426 0.754 0.756 0.803 0.942 0.788 0.727
Perplexity Filter 0.061 0.154 0.000 0.768 0.789 0.416 0.426 0.754 0.756 0.803 0.942 0.788 0.644
ICD 0.000 0.000 0.651 0.722 0.374 0.405 0.404 0.470 0.491 0.670 0.621 0.572 0.538
Paraphrase 0.061 0.154 0.145 0.625 0.458 0.377 0.287 0.426 0.435 0.645 0.682 0.546 0.463
Self-Reminder 0.000 0.000 0.362 0.634 0.353 0.382 0.383 0.407 0.395 0.451 0.472 0.484 0.432
SafeDecoding 0.000 0.015 0.000 0.000 0.044 0.051 0.065 0.052 0.038 0.035 0.067 0.070 0.042
Layer-AdvPatcher 0.034 0.146 0.754 0.696 0.516 0.369 0.408 0.463 0.476 0.704 0.813 0.693 0.589
MirrorShield (Ours) 0.000 0.010 0.000 0.000 0.048 0.021 0.036 0.013 0.015 0.014 0.021 0.054 0.022
@@
Layer-AdvPatcher 1.067 $\times$ 1.043 $\times$
MirrorShield(Ours) 1.058 $\times$ 1.064 $\times$
Table 2: This table summarizes ATGR of MirrorShield and the baseline defense approaches.
@@
Layer-AdvPatcher 64.7% 0.428 88.6% 0.519
MirrorShield(Ours) 68.6% 0.443 91.2% 0.535
@@
Metrics. Following previous work Xu et al. (2024), the attack success rate (ASR) is used to measure the effectiveness of the defense methods. The success of a jailbreak attack is evaluated by GPT-4o (Qi et al., 2023). The Average Token Generation Time Ratio (ATGR) is used to assess the time cost of all defense methods (Xu et al., 2024). Moreover, the WinRate and Rouge-L scores are used to evaluate the general performance of LLMs in dealing with harmless tasks (Jiang et al., 2024). Details on these metrics are deferred to Appendix B.
Implementation Details. The details of implementation settings are given in Appendix C.